OpenStudy (anonymous):

the town mayor plans to build a diagonal pathway for the rectangular town plaza whose length is 20 m longer than the width. if the pathway is 20 m shorter than twice the width, find the length of the pathway

OpenStudy (anonymous):



OpenStudy (ikram002p):

Hint :- phytagorian thm ^

OpenStudy (anonymous):

how?

OpenStudy (ikram002p):

\(\Huge\color{green}{c^2}=\color{blue}{a^2}+\color{gold }{b^2}\) 

OpenStudy (ikram002p):

\(\Huge (2x-20)^2 =x^2 +(x+20)^2\)

OpenStudy (ikram002p):

\(\large 4x^2-40x+400 =x^2 + x^2+20x+400 \)

OpenStudy (ikram002p):

then you would have :- \(\large 2x^2-60x =0 \) solve for x

OpenStudy (ikram002p):

\(\large x(2x-60)=0\) so two cases x=0 ( ignore ) 2x-60=0 \(\rightarrow x=30\)

OpenStudy (ikram002p):

the length of the pathway =(2x-20) sub x= 30 length = 2 × 30 -20 =60 -20 =40 m